# Dynamical and structural properties of an absorbing phase transition: a case study from granular systems

### v1

@@ V.4 Numerical results
In Fig. 10e, we give the evolution of the dynamic structure factor for the “$\Delta+\gamma$, strong damping” system as $k$ is varied. However, in this case, we know that the theory is completely wrong, partially because of the non-Gaussianity of the velocity distribution function but probably also due to additional effects not taken into account in the theory. We therefore normalize the dynamic structure factor by $\bm{S}_{\rho\rho}(k,0)$, this helps us eliminate in part the issue induced by a wrong prediction of the structure factor. At equilibrium, we could completely factorize the dynamic structure factor into a part independent on the noise used thanks to the fluctuation dissipation theorem. Indeed, the quantity $\bm{S}_{\rho\rho}(k,w)/\bm{S}_{\rho\rho}(k)$ would be purely computable from $\bm{M}$ without the need to include a noise term. In this non-equilibrium system, $\bm{S}_{\rho\rho}(k,0)$ is also related to $\bm{S}_{\rho T}(k)$ in a non-trivial way and the same argument cannot be made. This will however provide useful information. We see that, while the theory does not quantitatively predict the results from, the simulations, it captures well the change of region from II to I as $k$ is decreased. This leads us to conclude that the equilibrium noise we chose may be wrong, although the expressions for the transport coefficients are roughly correct.
@@
In light of Ref. 93, these results are quite surprising. In this study, a similar model was used where a constant energy $\Delta E$ is injected at collision instead of relying on a constant velocity injection $\Delta$ as in our case. Our systems are equivalent at the level of the density and momentum field but differ at the level of the temperature field due to different $G$:
@@
As clearly visible in these equations, our energy injection depends on the kinematic of the impact (and thus, on the temperature after the average on the collisions) while this is not the case in the system introduced in Ref. 93.
@@
Nonetheless, from Eq. (52), we also predict that the large $w$ tail of the dynamic structure factor is proportional to $\eta_{0}^{\parallel}$. But we see that this tail is very well predicted by the theory for the “$\Delta+\gamma$, weak damping” model in Fig. 10b, implying that from this observable we are using a correct value for the longitudinal viscosity. Moreover, we remark that according to the theory (Eq. (74) and (53)) both $\bm{S}_{\rho\rho}(k_{0},w\to 0)$ and $\bm{S}_{\rho\rho}(k_{0},w\to\infty)$ should be proportional to $\eta_{0}^{\parallel}$ which implies that, if the viscosity was the issue, we should observe discrepancies of similar order of magnitude at $w=0$ and $w\to\infty$, which is not the case.
@@
In this section, we only analyzed the coarse-grained system. A similar analysis could be performed for the realistic model. Analysis (not shown) indicate that the dynamic structure factor obtained in the realistic model is consistent with the picture given for the coarse-grained model, however, there are differences concerning the structure factor. In the realistic model, the structure factor never shows a strong $\bm{S}_{\rho\rho}(k)\sim k^{2}$ hyperuniformity. This is explained by the lack of conservation of center of mass of the noise. In an idealized quasi-2D system, the global damping is due to tangential friction happening at each grain-plate collision. Thus, the damping is applied stochastically and discretely. Compared to the idealized global damping of the coarse-grained model, this does not conserve the position of the center of mass and as was demonstrated in Ref. 197 forbids true hyperuniformity. In an experimental system, the roughness of the plate would also affect the dynamics and would be the most important effect. A particle would act closely to a Brownian particle and a global noise would be effectively applied [203, 204], forbidding hyperuniformity [197].
Finally, the dynamic structure factor of the system close to the transition, would merit further investigations. However, this is a topic for a separate study. In any case, the results are expected to diverge significantly from the predictions of a simple linear theory. Indeed, in the vicinity of the transition, even a Gaussian theory going beyond mean field is unable to predict the critical induced hyperuniformity [114]. Consequently, a renormalization group approach is necessary for a comprehensive understanding of the underlying physics [114].
@@ VI Conclusion
In conclusion, our investigation into the vibrofluidized granular model provides insights into the nature of absorbing phase transitions in systems with active collisions. We have demonstrated that by tuning parameters, it is possible to switch between continuous and discontinuous transitions, underscoring the critical role of system interactions. We find that the structural and dynamical properties of the active states close to the transitions are radically different according to the nature of the transition. We explain our findings through an equilibrium picture when possible, for example concerning a nucleation pathway to the absorbing state or an effective non-equilibrium thermodynamic potential. However, strongly non-equilibrium effects such as lack of Gaussianity or hyperuniformity were investigated and explained through kinetic theory and a fluctuating hydrodynamic description derived from the microscopic collision rule. Future work will focus on a more detailed experimental analysis to further validate the universality of our findings. Numerical investigations are still needed to better understand the tricritical point and the nucleation pathway in this system.
@@ Appendix A Argument and potential derivation of the overpopulated tail of the velocity distribution
We follow Refs. [154, 117, 116] to derivate an overpopulated tail in granular medium. We adapt their argument in order to potentially explain the overpopulated tail observed close to the continuous transition in Fig. 7. We note however that the physical mechanism leading to this overpopulated tail is different in the simple granular cooling and in our case.
@@
We proceed by noting that since the considered particle $i$ has a large velocity $v_{i}$, it will on average collide with particles of smaller velocity. Consequently, $\bm{v}_{ij}\simeq\bm{v}_{i}$ and $\bm{v}_{ij}\cdot\hat{\bm{\sigma}}_{ij}-2\Delta\simeq\bm{v}_{i}\cdot\hat{\bm{\sigma}}_{ij}$. For ease of theoretical computation and to deal with the post collisional velocities, with focus on 1D systems where we do not expect any differences for the velocity distribution in simulations. We further simplify the problem by setting $\alpha=1$ since the same overpopulated tail is also obtained in an elastic system [93]. These assumptions for positive $v_{i}$ lead to:
$-\gamma\dfrac{\partial}{\partial v_{i}}(v_{i}f)\simeq 2\chi\sigma(v_{i}f(v_{i})),$ (56)
@@
with $\beta=2\chi\sigma/\gamma$ and $Z$ an integration constant. Eq. (57) demonstrates the potential development of an exponential overpopulated tail. However, a rigorous numerical and theoretical analysis are necessary to validate our approximations. Notably, the assumption of homogeneity is uncontrolled, as we saw that the population of active particle has a drastic effect on the global temperature of the system. Moreover, 1D systems are known to exhibit peculiar thermalization [205, 206, 207, 208, 209]. For example, a system of 1D hard elastic particles would never reach equilibrium because the collision rule would simply exchange the velocities of the colliding particles. A Gaussian velocity distribution would never be reached and the initial distribution would be conserved. With our non-equilibrium collision rule, a 1D fluid can equilibrate. However, there might still be some differences compared to the $2d$ case, particularly regarding the efficiency of the energy spreading.
@@
Figure 11: $\bm{S}_{\rho\rho}(k)$, $\bm{S}_{u^{\parallel}u^{\parallel}}(k)$ and $\bm{S}_{u^{\perp}u^{\perp}}(k)$ as a function of $k$ close to the critical point. $N=10^{5}$, $\hat{\tau}\gamma=0.01$, $\alpha=0.95$, $\hat{\tau}\Delta/\sigma=0.015$ and $\phi=0.1512$.
@@ Appendix C Hyperuniform and clustered metastable fluids
Figure 12: a) Structure factor as a function of $k$ for our system with synchronization but with $\alpha=1$ at $\phi=0.22$, $N=10^{6}$ (very close to transition packing fraction) and $\phi=0.223$, $N=4\times 10^{5}$ (almost stable). b) Average waiting time for the metastable state ($\phi=0.22$) to reach the absorbing state as a function of the number of particle in the system. The average is performed over 20 independent runs. $\hat{\tau}\gamma=0.03$, $\hat{\tau}/\tau_{s}=0.13$ and $\hat{\tau}\Delta/\sigma=0.04$.
As discussed in the main text, we did not observe a clear hyperuniform scaling in the “metastable” fluid phase, nor its stabilization with increasing system size $N$, as reported in Ref. 103. This is unexpected given the model similarities. However, in contrast to Ref. 103, where dissipation arises solely from viscous drag, our model with $\alpha<1$ includes dissipative collisions, which are responsible for driving the system to the absorbing state. The purely dissipative collisions between synchronized particles cause clustering, therefore inducing a peak at a wavevector $k^{*}$ in the structure factor of the metastable fluid. These clusters appear to aid the system in reaching an absorbing state via a mechanism akin to equilibrium nucleation. Indeed, the system size $L^{*}$ at which a sharp drop in waiting time is found in Fig. 6 corresponds roughly to $2\pi/k^{*}$. However, it remains highly unclear why such phenomena facilitate the transition.
@@
We now turn our analysis to a metastable fluid with a discontinuous transition but with $\alpha=1$. In Fig. 12a), we present the structure factor in the metastable regime (close to the discontinuous transition) in our model, but with $\alpha=1$ where collisions are elastic, and therefore clustering does not occur. In contrast to the metastable fluid with $\alpha\neq 1$ (Fig. 7), we do not observe any peak in the structure factor. Instead, a hyperuniform scaling close to $k^{1.2}$ (but not incompatible with $k^{2}$ scaling at lower $k$) is visible for the metastable fluid ($\phi=0.22$), in agreement with results from Ref. 103. We also note that a very long $k^{0.3}$ scaling is observed close to the metastable fluid. Additionally, the average waiting time $\langle t_{wait}\rangle$ increases monotonically with system size (or $N$) as given in Fig. 12b). This is in contrast with the non-monotonic behavior observed in Fig. 6. Thus, when dissipative collisions are absent, our system behaves similarly to the one studied in Ref. 103, allowing for a hyperuniform scaling to emerge in the metastable fluid. This scaling seemingly prohibits nucleation in the limit $N\to\infty$.
@@ Appendix D Tricritical point and continuous transition with synchronization
Figure 13: a) Transition from the discontinuous to continuous transition with synchronization as $\alpha$ is decreased: $N=10000$, $\hat{\tau}\Delta/\sigma=0.025$, $\hat{\tau}/\tau_{s}=0.25$ and $\hat{\tau}\gamma=0.01$. b) Is a close-up on the transition for $\alpha=0.8$ with a scaling analysis showing that the last active steady state $\Delta T_{ss}$ goes to 0 with system size, proving the continuity of the transition. The dashed line is the expected power law with the exponent from the conserved directed percolation universality class.
As discussed in the main text, even with synchronization, we can observe a continuous transition because the synchronization time competes with a finite mean collision time in the active region. Thus, at a large but finite synchronization time $\tau_{s}$, significantly longer than the minimum mean collision time in active regions, synchronization will not influence the transition, which remains continuous. A less straightforward example of a continuous transition at finite $\tau_{s}$ occurs when $\alpha\ll 1$ as illustrated by Fig. 13a. Indeed, we observe that by decreasing $\alpha$, the transition becomes continuous. The continuity of the transition is made explicit in Fig. 13b where a scale analysis is performed for $\alpha=0.8$ and shows that the temperature of the last active steady state tends to approach 0 with increasing system size. While the theory cannot predict quantitatively this transition, it still gives a qualitative answer. As evidenced from Eq. (29), in the case of small $\alpha$, a discontinuous transition is theoretically predicted to occur at a small $T_{a}$. However, the population of inactive particles will reach 0 before the theoretical discontinuous transition happens, leading to a continuous transition instead. Similar transitions have been investigated in sandpiles belonging to the conserved directed percolation universality class [51]. In those systems, the mean field approximation has been shown to quantitatively fail considerably around the passage from the discontinuous to continuous transition, notably, it has been found that the continuous transition for a range of parameters, is fluctuation induced.
This behavior implies the existence of a non-equilibrium tricritical point in our system. Such a phenomenon has been documented and thoroughly investigated in a closely related system in Ref. 102. It would be interesting to observe a tricritical point in the realistic system as well. It will require further exploration, for example through the use of very inelastic beads such as plastic ones. Note that a tricritical point as been observed as well for the liquid-solid phase transition of a quasi-2D granular gas [210, 211, 212].

### v2

@@ V.4 Numerical results
In Fig. 10e, we give the evolution of the dynamic structure factor for the $\Delta+\gamma$, strong damping” system as $k$ is varied. However, in this case, we know that the theory is completely wrong, partially because of the non-Gaussianity of the velocity distribution function but probably also due to additional effects not taken into account in the theory. We therefore normalize the dynamic structure factor by $\bm{S}_{\rho\rho}(k,0)$, this helps us eliminate in part the issue induced by a wrong prediction of the structure factor. At equilibrium, we could completely factorize the dynamic structure factor into a part independent on the noise used thanks to the fluctuation dissipation theorem. Indeed, the quantity $\bm{S}_{\rho\rho}(k,w)/\bm{S}_{\rho\rho}(k)$ would be purely computable from $\bm{M}$ without the need to include a noise term. In this non-equilibrium system, $\bm{S}_{\rho\rho}(k,0)$ is also related to $\bm{S}_{\rho T}(k)$ in a non-trivial way and the same argument cannot be made. This will however provide useful information. We see that, while the theory does not quantitatively predict the results from, the simulations, it captures well the change of region from II to I as $k$ is decreased. This leads us to conclude that the equilibrium noise we chose may be wrong, although the expressions for the transport coefficients are roughly correct.
@@
In light of Ref. lei2019hydrodynamics, these results are quite surprising. In this study, a similar model was used where a constant energy $\Delta E$ is injected at collision instead of relying on a constant velocity injection $\Delta$ as in our case. Our systems are equivalent at the level of the density and momentum field but differ at the level of the temperature field due to different $G$:
@@
As clearly visible in these equations, our energy injection depends on the kinematic of the impact (and thus, on the temperature after the average on the collisions) while this is not the case in the system introduced in Ref. lei2019hydrodynamics.
@@
Nonetheless, from Eq. (52), we also predict that the large $w$ tail of the dynamic structure factor is proportional to $\eta_{0}^{\parallel}$. But we see that this tail is very well predicted by the theory for the $\Delta+\gamma$, weak damping” model in Fig. 10b, implying that from this observable we are using a correct value for the longitudinal viscosity. Moreover, we remark that according to the theory (Eq. (74) and (53)) both $\bm{S}_{\rho\rho}(k_{0},w\to 0)$ and $\bm{S}_{\rho\rho}(k_{0},w\to\infty)$ should be proportional to $\eta_{0}^{\parallel}$ which implies that, if the viscosity was the issue, we should observe discrepancies of similar order of magnitude at $w=0$ and $w\to\infty$, which is not the case.
@@
In this section, we only analyzed the 2D system. A similar analysis could be performed for the realistic quasi-2D model. Analysis (not shown) indicate that the dynamic structure factor obtained in the realistic quasi-2D model is consistent with the picture given for the effective 2D model, however, there are differences concerning the structure factor. In the realistic quasi-2D model, the structure factor never shows a strong $\bm{S}_{\rho\rho}(k)\sim k^{2}$ hyperuniformity. This is explained by the lack of conservation of center of mass of the noise. In an idealized quasi-2D system, the global damping is due to tangential friction happening at each grain-plate collision. Thus, the damping is applied stochastically and discretely. Compared to the idealized global damping of the effective 2D model, this does not conserve the position of the center of mass and as was demonstrated in Ref. maire2024enhancing forbids true hyperuniformity. In an experimental system, the roughness of the plate would also affect the dynamics and would be the most important effect. A particle would act closely to a Brownian particle and a global noise would be effectively applied [puglisi2012structure, sarracino2010granular], forbidding hyperuniformity [maire2024enhancing].
Finally, the dynamic structure factor of the system close to the transition, would merit further investigations. However, this is a topic for a separate study. In any case, the results are expected to diverge significantly from the predictions of a simple linear theory. Indeed, in the vicinity of the transition, even a Gaussian theory going beyond mean field is unable to predict the critical induced hyperuniformity [ma2023theory]. Consequently, a renormalization group approach is necessary for a comprehensive understanding of the underlying physics [ma2023theory].
@@ VI Conclusion
In this work, we conducted an in-depth investigation of the system introduced in Ref. maire2024interplay. We demonstrated that an absorbing phase transition can occur in a granular system, with the nature of the transition determined by the presence or absence of synchronization. Building upon our previous findings, we analyzed in greater detail the structural properties of the system, both near the transition and within the active phase. We found that the critical point of the continuous transition exhibits a highly heterogeneous steady state, characterized by avalanche dynamics and non-trivial correlations, with critical exponents consistent with those of the conserved directed percolation universality class. In contrast, the discontinuous transition displays a more homogeneous state up to the onset of nucleation, which subsequently drives the system into a cooling regime reminiscent of free granular cooling and its associated instabilities. Furthermore, the active state of the corresponding two-dimensional effective model was shown to possess long-range correlations and to display hyperuniformity. Building on the theoretical framework established in Ref. maire2024interplay, we derived from the microscopic degrees of freedom, the field theory of the conserved directed percolation for our particle system, enabling a quantitative prediction of the homogeneous, asynchronized state close and far from the transition. By incorporating synchronization at the mean-field level, we also quantitatively captured the discontinuous transition. Finally, by including fluctuations around the homogeneous state, our extended theory successfully accounts for all observed correlation functions of the non-equilibrium liquid, as well as its emergent hyperuniformity.
Future work will focus on a more detailed experimental analysis to further validate the universality of our findings. Numerical investigations are still needed to better understand the tricritical point and the nucleation pathway in this system.
@@ Appendix A Argument and potential derivation of the overpopulated tail of the velocity distribution
We follow Refs. [brilliantov2004kinetic, van1998velocity, esipov1997granular] to derivate an overpopulated tail in granular medium. We adapt their argument in order to potentially explain the overpopulated tail observed close to the continuous transition in Fig. 7. We note however that the physical mechanism leading to this overpopulated tail is different in the simple granular cooling and in our case.
@@
We proceed by noting that since the considered particle $i$ has a large velocity $v_{i}$, it will on average collide with particles of smaller velocity. Consequently, $\bm{v}_{ij}\simeq\bm{v}_{i}$ and $\bm{v}_{ij}\cdot\hat{\bm{\sigma}}_{ij}-2\Delta\simeq\bm{v}_{i}\cdot\hat{\bm{\sigma}}_{ij}$. For ease of theoretical computation and to deal with the post collisional velocities, with focus on 1D systems where we do not expect any differences for the velocity distribution in simulations. We further simplify the problem by setting $\alpha=1$ since the same overpopulated tail is also obtained in an elastic system [lei2019hydrodynamics]. These assumptions for positive $v_{i}$ lead to:
$\gamma\dfrac{\partial}{\partial v_{i}}(v_{i}f)\simeq 2\chi\sigma(v_{i}f(v_{i})),$ (56)
@@
with $\beta=2\chi\sigma/\gamma$ and $Z$ an integration constant. Eq. (57) demonstrates the potential development of an exponential overpopulated tail. However, a rigorous numerical and theoretical analysis are necessary to validate our approximations. Notably, the assumption of homogeneity is uncontrolled, as we saw that the population of active particle has a drastic effect on the global temperature of the system. Moreover, 1D systems are known to exhibit peculiar thermalization [spohn2014nonlinear, spohn2016fluctuating, eltohfa2024simulations, movsko1995thermalization, onorato2023wave]. For example, a system of 1D hard elastic particles would never reach equilibrium because the collision rule would simply exchange the velocities of the colliding particles. A Gaussian velocity distribution would never be reached and the initial distribution would be conserved. With our non-equilibrium collision rule, a 1D fluid can equilibrate. However, there might still be some differences compared to the $2D$ case, particularly regarding the efficiency of the energy spreading.
@@
Figure 11: $\bm{S}_{\rho\rho}(k)$, $\bm{S}_{u^{\parallel}u^{\parallel}}(k)$ and $\bm{S}_{u^{\perp}u^{\perp}}(k)$ as a function of $k$ close to the critical point. $N=10^{5}$, $\alpha=0.95$, $\Delta/(\sigma\gamma)=1.5$ and $\phi=0.1512$.
@@ Appendix C Hyperuniform and clustered metastable fluids
Figure 12: a) Structure factor as a function of $k$ for our system with synchronization but with $\alpha=1$ at $\phi=0.22$, $N=10^{6}$ (very close to transition packing fraction) and $\phi=0.223$, $N=4\times 10^{5}$ (almost stable). b) Average waiting time for the metastable state ($\phi=0.22$) to reach the absorbing state as a function of the number of particle in the system. The average is performed over 20 independent runs. $1/(\gamma\tau_{s})=4.33$ and $\Delta/(\gamma\sigma)=1.33$.
The model presented in Ref. [lei2021barrier, lei2023does] is similar to ours. While it does not contain synchronization, the collision dynamics is effectively similar to ours. In their model, if two particles collide with an energy smaller than some threshold, then the particles undergo elastic collision, otherwise they gain a finite amount of energy. A viscous drag is also applied during the free flight. In our model, particles that do not have a lot of energy are, on average, those which traveled for a long time and are slowed down due to drag. Hence, they are likely to be synchronized and will therefore undergo collisions with $\Delta=0$. This might mimic the effect of the energy-activated active collision of the models studied by Lei et al. Thus, the rough phenomenology of both models is similar even though Refs. lei2021barrier, lei2023does do not consider dissipative collisions, and drag is de facto their only source of dissipation. Interestingly, they find that the pathway of the metastable fluid to the absorbing state is seemingly not driven by a nucleation process but by large wavelength fluctuations which are suppressed due to the non-equilibrium nature of the system. Indeed, they find that the metastable fluid has suppressed density fluctuations. On this basis, they argue that the non-equilibrium equivalent of the equilibrium metastable state becomes kinetically stable in the thermodynamic limit, due to the suppression of large wavelength fluctuations. The shrinking of the metastable region with system size is also observed using the minimal field theory of the conserved directed percolation universality class, modified to obtain a first order transition [di2016self, PhysRevResearch.2.013318]. This is at odds with our results, as well as what was found in Ref. neel2014dynamics. We believe the discrepancy is caused by the presence of dissipative collisions in our system that induces the formation of clusters, which enhance long wavelength fluctuations and allow for the regular nucleation pathway to the transition. Indeed, in contrast to Ref. lei2023does, where dissipation arises solely from viscous drag, our model with $\alpha<1$ includes dissipative collisions, which are responsible for driving the system to the absorbing state. The purely dissipative collisions between synchronized particles cause clustering, therefore inducing a peak at a wavevector $k^{*}$ in the structure factor of the metastable fluid. These clusters appear to aid the system in reaching an absorbing state via a mechanism akin to equilibrium nucleation. Indeed, the system size $L^{*}$ at which a sharp drop in waiting time is found in Fig. 6 corresponds roughly to $2\pi/k^{*}$. However, it remains highly unclear why such phenomena facilitate the transition.
@@
We now turn our analysis to a metastable fluid with a discontinuous transition but with $\alpha=1$. In Fig. 12a), we present the structure factor in the metastable regime (close to the discontinuous transition) in our model, but with $\alpha=1$ where collisions are elastic, and therefore clustering does not occur. In contrast to the metastable fluid with $\alpha\neq 1$ (Fig. 7), we do not observe any peak in the structure factor. Instead, a hyperuniform scaling close to $k^{1.2}$ (but not incompatible with $k^{2}$ scaling at lower $k$) is visible for the metastable fluid ($\phi=0.22$), in agreement with results from Ref. lei2023does. We also note that a very long $k^{0.3}$ scaling is observed close to the metastable fluid. Additionally, the average waiting time $\langle t_{wait}\rangle$ increases monotonically with system size (or $N$) as given in Fig. 12b). This is in contrast with the non-monotonic behavior observed in Fig. 6. Thus, when dissipative collisions are absent, our system behaves similarly to the one studied in Ref. lei2023does, allowing for a hyperuniform scaling to emerge in the metastable fluid. This scaling seemingly prohibits nucleation in the limit $N\to\infty$.
@@ Appendix D Tricritical point and continuous transition with synchronization
Figure 13: a) Transition from the discontinuous to continuous transition with synchronization as $\alpha$ is decreased: $N=10000$, $\Delta/(\gamma\sigma)=2.5$ and $1/\gamma\tau_{s}=25$. b) Is a close-up on the transition for $\alpha=0.8$ with a scaling analysis showing that the last active steady state $\Delta T_{ss}$ goes to 0 with system size, proving the continuity of the transition. The dashed line is the expected power law with the exponent from the conserved directed percolation universality class.
As discussed in the main text, even with synchronization, we can observe a continuous transition because the synchronization time competes with a finite mean collision time in the active region. Thus, at a large but finite synchronization time $\tau_{s}$, significantly longer than the minimum mean collision time in active regions, synchronization will not influence the transition, which remains continuous. A less straightforward example of a continuous transition at finite $\tau_{s}$ occurs when $\alpha\ll 1$ as illustrated by Fig. 13a. Indeed, we observe that by decreasing $\alpha$, the transition becomes continuous. The continuity of the transition is made explicit in Fig. 13b where a scale analysis is performed for $\alpha=0.8$ and shows that the temperature of the last active steady state tends to approach 0 with increasing system size. While the theory cannot predict quantitatively this transition, it still gives a qualitative answer. As evidenced from Eq. (29), in the case of small $\alpha$, a discontinuous transition is theoretically predicted to occur at a small $T_{a}$. However, the population of inactive particles will reach 0 before the theoretical discontinuous transition happens, leading to a continuous transition instead. Similar transitions have been investigated in sandpiles belonging to the conserved directed percolation universality class [di2016self]. In those systems, the mean field approximation has been shown to quantitatively fail considerably around the passage from the discontinuous to continuous transition, notably, it has been found that the continuous transition for a range of parameters, is fluctuation induced.
This behavior implies the existence of a non-equilibrium tricritical point in our system. Such a phenomenon has been documented and thoroughly investigated in a closely related system in Ref. lei2021barrier. It would be interesting to observe a tricritical point in the realistic system as well. It will require further exploration, for example through the use of very inelastic beads such as plastic ones. Note that a tricritical point as been observed as well for the liquid-solid phase transition of a quasi-2D granular gas [guzman2018critical, castillo2012fluctuations, castillo2015universality].